OpenStudy (anonymous):

x^2 + y^2 = 4, find d^2y/dx^2

OpenStudy (anonymous):

i can find the first derivative, -x/y and i know the answer should be 4/y^3

OpenStudy (turingtest):

are you using implicit differentiation yet?

OpenStudy (anonymous):

yeah thats exactly it, sorry i can find differnetiations and i know i'm supposed to differntiate my differentiation but it won't come out properly

OpenStudy (anonymous):

dy/dx -> 2x + 2y * (dy/dx) = 0 dy /dx = -x/y d^2y / dx^2 -> 2 + 2 * (dy/dx * dy/dx + y * (d^2y/ dx^2)) =0 2 + 2 * ((dy/dx)^2 + y * (d^2y/dx^2)) = 0 2 + 2 * (x^2/y^2 + y * d^2y/dx^2) =0 -(1 + x^2/y^2) / y = d^2y/dx^2 -(y^2 + x^2) / y^3 y^2 = 4-x^2 so : -(4-x^2 + x^2) / y^3 = -4/y^3